OpenStudy (anonymous):

Find the next three terms of the sequence –12, 24, –48, 96, . . . A.–192, 384, –768 B.96, –192, 384 C.192, –384, 768 D.132, –156, 132

OpenStudy (anonymous):

12 12*2 = 24 24*2 = 48 48*2 = 96 Continue that pattern. And the signs alternate.

OpenStudy (anonymous):

A